# When More is Less: Understanding Chain-of-Thought Length in LLMs

### v1

@@ 2 Related Work
Chain of Thought Large Language Models (LLMs) (Brown et al., 2020) have demonstrated remarkable abilities in complex reasoning tasks by breaking down challenging problems into intermediate steps before arriving at the final answer (Wei et al., 2022). Numerous researchers have proposed various approaches to enhance the CoT reasoning capabilities of LLMs. Least-to-most prompting (Zhou et al., 2023) decomposes a complex problem into a series of simpler sub-problems, solving them sequentially, where the solution to each subproblem builds upon the answers to previously solved sub-problems. Tree of thoughts (Yao et al., 2023) enables LLMs to engage in deliberate decision-making by exploring multiple reasoning paths, self-evaluating options, and dynamically adjusting the reasoning process through backtracking or look-ahead strategies to make globally optimal choices. Similarly, Divide-and-Conquer methods (Zhang et al., 2024; Meng et al., 2024) divide the input sequence into multiple sub-inputs, which can significantly improve LLM performance in specific tasks. Despite their differences, these methods share a common characteristic: they all treat the CoT process as a framework for decomposition and subtask-solving. Similarly, our study adopts this perspective.
CoT Understanding In addition to the methods mentioned above, many works aim to formalize the CoT process and explore why it is effective. Circuit complexity theory has been used to analyze the computational complexity of problems that transformers can solve with and without CoT, providing a theoretical understanding of CoT’s effectiveness (Feng et al., 2023; Li et al., 2024b). Cui et al. (2024) theoretically demonstrate that, compared to Stepwise ICL, integrating reasoning from earlier steps (Coherent CoT) enhances transformers’ error correction capabilities and prediction accuracy. Ton et al. (2024) quantify the information gain at each reasoning step in an information-theoretic perspective to understand the CoT process. Furthermore, Li et al. (2024a) show that fast thinking without CoT results in larger gradients and greater gradient differences across layers compared to slow thinking with detailed CoT, highlighting the improved learning stability provided by the latter. Ye et al. (2024) investigate CoT in a controlled setting by training GPT-2 models on a synthetic GSM dataset, revealing hidden mechanisms through which language models solve mathematical problems. Unlike these theoretical studies, our work focuses on the impact of different lengths of CoT on final performance and tries to understand CoT from task decomposition and error accumulation perspective.
Overthinking With the remarkable success of OpenAI’s o1 model, test-time computation scaling has become increasingly important. More and more works (Snell et al., 2024; Chen et al., 2024d; Wu et al., 2024; Brown et al., 2024) have explored the scaling laws during inference using various methods, such as greedy search, majority voting, best-of-n, and their combinations. They concluded that with a compute-optimal strategy, a smaller base model can achieve non-trivial success rates, and test-time compute can outperform larger models. This highlights the importance of designing optimal inference strategies.
However, Chen et al. (2024a) hold a contrastive opinion that in some cases, the performance of the Best-of-N method may decline as $N$ increases. Similarly, the overthinking phenomenon (Chen et al., 2024c) becomes more and more important as o1-like reasoning models allocate excessive computational resources to simple problems (e.g., $2+3=5$) with minimal gains. These findings indicate the need to balance computation based on model capabilities and task difficulty. In our study, we focus on different types of CoT reasoning, categorized by CoT length. Moreover, we theoretically identify a balanced CoT strategy that adapts to model size and task difficulty, optimizing performance under these constraints.
3 Influence of Chain-of-Thought Length on Arithmetic Tasks
@@
To begin, we aim to empirically investigate the relationship between reasoning performance and CoT length. Therefore, we need to control a given model to generate reasoning chains of varying lengths for a specific task. Unfortunately, no existing real-world dataset or model fully meets these strict requirements. Real-world reasoning tasks, such as GSM8K or MATH (Cobbe et al., 2021; Hendrycks et al., 2021), do not provide multiple solution paths of different lengths, and manually constructing such variations is challenging. Moreover, it is difficult to enforce a real-world model to generate a diverse range of reasoning paths for a given question.
Given these limitations, we begin our study with experiments on synthetic datasets. Notably, even when working with real-world datasets, we observe behavioral patterns that align well with the findings derived from synthetic data (Section 5.1).
3.1 Problem Formulation
To investigate the effect of CoT length in a controlled manner, we design a synthetic dataset of simplified arithmetic tasks with varying numbers of reasoning steps in the CoT solutions. The necessity and rationale for simplified arithmetic tasks will be further discussed in the Appendix B.
Definition 3.1 (Problem).
@@ 3.1 Problem Formulation
To ensure that CoT solutions of the same length have equal difficulty for a specific problem, we assume that each reasoning step performs the same operations within a single CoT process. A more rigorous discussion will be conducted in Appendix B.2.
Definition 3.2 (Solution).
@@
3.2 Experiment Setup
(a) Small model with 6 layers
(b) Large model with 9 layers
Figure 3: The envelope curve shows how the performance of optimal operators in a single step ($t$) changes as the task becomes more difficult. The different colors of the envelope curve correspond to the best-performing operators in one step ($t$) for the given set of total operators.
Dataset The datasets used for training different models are identical and are constructed from tasks with varying total operators $T$ and solutions with different $t$.
Model Since model depth plays a critical role in mathematical reasoning (Ye et al., 2024), we trained GPT-2 models with varying numbers of layers, keeping all other parameters (e.g., heads, dimensions) constant, to assess the impact of model capacity. Our experimental results provide evidence that models with more layers are capable of performing more operations in a single step, without the need for CoT.
Train and Test For each problem, we train the model to start generating from the control token <t> to ensure that it can independently determine which CoT solution to use. During testing, we guide the model to produce the required $t$-hop CoT by inserting the control token <t> right before the question. More details can be found in Appendix D.
3.3 Experimental Results
U-curve For convenience, we present how the final accuracy changes as the number of operators performed per step $t$ increases, which corresponds to a decrease in the number of reasoning steps, for tasks of varying difficulty (total operators). Figures 1(a) and 1(b) show the performance of small and large models on easy tasks (16, 24, 32 operators) and hard tasks (48, 56, 64 operators) respectively.
The results (Figure 1) align well with our theory, demonstrating that as the number of reasoning steps increases, the final performance initially improves and then declines. Subsequent experiments will explore how the optimal CoT length changes with model capability and task difficulty.
Envelope Curve Figure 3(a) and 3(b) illustrate the envelope curve, where different colors represent the optimal single-step length. The results indicate that as the task becomes more challenging, CoT with a larger single-step reasoning length $t$ achieves the best performance. This can be interpreted as a mechanism to regulate the total number of CoT steps—shorter single-step lengths require more CoT steps to complete the task.
Figure 4: Heat map of optimal CoT length in different model sizes and task difficulties.
Optimal CoT length shifts To further investigate how the optimal CoT length changes with model capability and task difficulty, we trained models of varying sizes (ranging from 5 to 9 layers) on tasks of different difficulties (from 16 to 64 operators). For each combination of model size and task difficulty, we recorded the optimal number of reasoning steps.
Figure 4 illustrates two key findings. First, an increase in the number of reasoning steps is beneficial for solving more challenging problems, indicating that harder tasks require more steps to achieve optimal performance. Second, the optimal number of reasoning steps decreases as the model size increases, suggesting that stronger models can handle more complex reasoning within fewer steps.
4 Theoretical Analysis
In this section, we provide a theoretical analysis of the CoT process for the simplified arithmetic tasks defined above and explain the empirical results observed in synthetic datasets. All proofs of the paper are deferred to Appendix F.
4.1 Setup
Let $N\in\mathbb{N^{+}}$ represent the total number of steps in the CoT process. As defined earlier, $T$ denotes the total number of operators in the given question, and $t=\left\lceil\frac{T}{N}\right\rceil$ represents the number of operators processed in each reasoning step. We use $t_{i}$ to denote the subtask in the $i$-th reasoning step (e.g., $2+1$ in Eq. (2)), and $a_{i}$ to represent the corresponding answer (e.g., $3$ in Eq. (2)).
Definition 4.1.
Given task $q$ with total operators $T$ (Definition 3.1) and model $\theta$, to a specific $N$, we define an $N$ step ($t$-hop in Definition 3.2) CoT process as
$P(a_{N}|q,\theta)=\prod^{N}_{i=1}P(t_{i}|H_{i-1},q,\theta)P(a_{i}|t_{i},H_{i-1},q,\theta),$
where $H_{k}=[t_{1},a_{1},\cdots,t_{k},a_{k}]$ collects the first $k$ steps in the $N$-step CoT process, and $a_{N}$ is not only the answer of the final subtask $t_{i}$, but the answer to the whole task $q$ as well.
Let $a_{i}^{*}$ denote the correct answer to subtask $t_{i}$, and $t_{i}^{*}$ the correct next subtask given the history reasoning steps $H_{i-1}$, total task $q$, and total step number $N$. To estimate the final accuracy $A(N)=P(a_{N}=a_{N}^{*}|q,\theta)$, we need to separately estimate the sub-question accuracy $P(t_{i}=t_{i}^{*}|H_{i-1},q,\theta)$ and the sub-answer accuracy $P(a_{i}=a^{*}|t_{i},H_{i-1},q,\theta)$.
For the sub-question accuracy, as observed in our experiments, the loss for tokens appearing in $t_{i}$ remains almost constant across different values of $N$ and model sizes (see Appendix C). We assume that the noise affecting tokens in each subtask $t_{i}$, denoted as $\sigma(T)\in[0,1)$, is positively correlated with the total number of operators $T$. Intuitively, as the number of operators increases, extracting the correct subtask becomes more challenging.
For the sub-answer accuracy, it is clear that when given subtask $t_{i}$, $P(a_{i}=a^{*}|t_{i},H_{i-1},q,\theta)$ is independent of the history reasoning steps $H_{i-1}$ and is only influenced by the model $\theta$ and the difficulty of the subtask $t_{i}$. For each model, we define its capability $M$ based on the reasoning boundary (Chen et al., 2024b). Specifically, we consider $M$ as the maximum number of operators the model can compute in a single step without CoT.
$M=M(\theta)=\max_{t}\left\{P(a_{i}=a^{*}|t_{i},\theta)>0,|t_{i}|=t\right\},$ (4)
where $|t_{i}|$ refers to the number of operators in subtask $t_{i}$. In the following discussion, we focus only on CoT chains that do not exceed the model’s capability, which is $t<M.$ Hence, we define the error rate of each subtask answer as $E(N,M,T)\in[0,1)$. {restatable}propositionfinalacc The total accuracy of $N$-step reasoning is
$A(N)=\alpha\left((1-E(N,M,T))(1-\sigma(T))\right)^{N},$ (5)
where $\alpha$ denotes a constant value independent of $N$.
Proposition 5 establishes a differentiable functional relationship between total accuracy and CoT length $N$. Once we obtain estimates for $E(N,M,T)$ and $\sigma(T)$, we can determine the optimal $N(M,T)$. For simplicity, in the following discussion, we allow $N$, $M$, and $T$ to be real numbers. In Section 4.2, we analyze the case of a linear error rate, while in Section 4.3, we explore more general error functions.
4.2 A Simple Case with Linear Error
An estimation of $\sigma(T)$. To simplify the setting, we assume $\sigma(T)=\frac{T}{C}$, where $C$ is a hyperparameter representing the maximum number of operators the model can handle, which is solely influenced by the training data. To ensure that the model is capable of generating a reasonable subtask (even if it is incorrect), we consider a finite range $T\in[0,0.9C]$, ensuring that the subtask accuracy rate $1-\sigma(T)$ remains within $[0.1,1]$.
An estimation of $E(N,M,T)$. We define the model answer’s error rate $E(N,M,T)=\frac{T}{N}/{M}$ as the ratio between the number of subtask operators and the model’s capacity $M$ (Eq. (4)) that the maximum number of operators the model can compute in a single step. Therefore, $01-\frac{T}{NM}>$ 0.
According to Proposition 5, a simplified total accuracy of $N$-step reasoning is
$A(N)=\alpha\left(\left(1-\frac{T}{C}\right)\left(1-\frac{T}{NM}\right)\right)^{N}.$ (6)
{restatable}
theoremOptimalN In a simplified setting, for a given model capability $M$ and task difficulty $T$, the final accuracy $A(N)$ (Eq. (6)) increases initially and then decreases as the number of reasoning steps $N$ increases. Besides, there exists an optimal
$N(M,T)=\frac{TZ}{M(Z+1)}$ (7)
that maximizes the final accuracy, where $Z=W_{-1}\left(-\frac{1-\frac{T}{C}}{e}\right)$, and $W_{-1}(x)$ denotes the smaller branch of the Lambert function, which satisfies the equation $we^{w}=x$.
Theorem 4.2 shows the optimal $N(M,T)$ is only affected by two factors $M$ and $T$. Through the expression of $N$, we can naturally derive the following three observations about the trend of $N$ with respect to changes in $M$ or $T$:
Corollary 4.2.
$\frac{T}{N(T,M)}=M\left(1+\frac{1}{Z}\right)$ increases monotonically with $T$, which aligns the envelope curve result well.
{restatable}
corollaryNwithT $N(M,T)$ increases monotonically with $T$, which shows that as the total task becomes more difficult, the optimal number of reasoning steps increases.
Corollary 4.3.
$N(M,T)$ decreases monotonically with $M$, which shows that as the capability of the model becomes stronger, it is easier for the model to solve one-step subtask reasoning, which leads to the decrease of the total task decomposition number.
4.3 Extension to General Error Functions
In the simple case above, we discussed the trend of overall accuracy with respect to $N$ and the variation of optimal $N$ with $M$ and $T$, assuming the subtask error rate is a linear function. In the following discussion, we aim to derive conclusions corresponding to more general error rate functions. We find that as long as the error function satisfies some basic assumptions, the above conclusions still hold. A detailed discussion of basic assumptions will be conducted in Appendix E.
{restatable}
theoremgeneral For any noise function $10<\sigma(T)<$ 1 and a subtask error rate function $10<E(N,M,T)<$ 1 satisfying Assumption E.1 and E.2, a general final accuracy function $A(N)$ (Proposition 5) has the following properties:
  • $\lim_{N\rightarrow+\infty}A(N)\rightarrow 0$
  • If $A(N)$ has the point of maximum value $N^{*}>1$, then $N^{*}$ has a lower bound
    $N^{*}\geq N(M,T)=E^{-1}\left(1-\frac{1}{e^{2}(1-\sigma(T))}\right),$ (8)
where $E^{-1}$ is the inverse function of $E(N,M,T)$ with respect to $N$, which has the same monotonicity as $E$ with respect to $M$ and $T$. Therefore, we have similar corollaries as in the linear case.
Corollary 4.4.
As the model becomes stronger, $E^{-1}$ decreases monotonically with respect to $M$, which leads to decrease of $N(M,T)$.
Corollary 4.5.
As the task becomes harder, $E^{-1}$ is monotonically increasing with respect to $T$, which leads to an increase of $N(M,T)$.
5 Empirical Examination of Optimal CoT Length
Following the theoretical analyses of optimal CoT length in Section 4, we further validate these insights on both real-world LLM behaviors at test time and its influence on CoT training.
5.1 Optimal CoT Length of LLMs on MATH
(a) Optimal length v.s. Longest length
(b) Optimal length v.s. Model size
(c) Optimal length v.s. Task difficulty
Figure 5: Evaluation on MATH datasets with Qwen2.5 Series Instruct models.
To validate our theory on real-world tasks with LLMs, we consider the MATH (Hendrycks et al., 2021) algebra dataset (Level 5), which includes challenging competition-level mathematics problems requiring multi-step reasoning. We select Qwen2.5 series Instruct models (Qwen et al., 2025) to investigate behaviors among models with different capabilities. More details can be found in Appendix A.1.
Optimal Length and Model Capability In this section, we randomly select 30 questions from the dataset, generating 60 samples for each question, resulting in a total of 1,800 samples for each model. The results, shown in Figure 5(a), demonstrate that for each model, the longest number of steps is not the best one. Additionally, the optimal CoT length decreases as the model size increases, from 14 steps for the 1.5B model to 4 steps for the 72B model (Figure 5(b)). This trend aligns with our theory that larger models perform better with fewer reasoning steps, as they are more capable of effectively handling single-step reasoning.
Impact of Task Difficulty In this part, we evaluate how the optimal CoT length changes with task difficulty. To achieve this, we randomly select 100 questions from the dataset, generating 60 samples for each question to ensure the number of reasoning steps is calculated accurately. We assume that a question with lower accuracy among the 60 samples is more difficult for the model, using 1-accuracy as a proxy for the relative difficulty of each question(the larger the harder).
We then plot a scatter plot of accuracy versus optimal CoT length and calculate the correlation. As shown in Figure 5(c), a significant correlation ($p=1e-8<<0.05$) between task difficulty and optimal CoT length. Results on different models are shown in Appendix A.2. This finding supports our theory that harder tasks require more steps to solve effectively.
5.2 Implications for Training with CoT Data
In the previous synthetic dataset experiments (Section 3), we generated solutions with varying step lengths by training our model on data sampled with random CoT lengths. Now that we have identified the optimal CoT length, an important question arises:
Can we construct a dataset that contains only CoT solutions with optimal lengths, tailored to the current model size and task difficulty?
To explore this, we conduct experiments on a synthetic dataset. The baseline model is trained on CoT step lengths following a uniform distribution, while the optimal model is trained exclusively on the optimal CoT lengths identified in Figure 4.
Figure 6: Comparison of training on CoTs with optimal lengths versus random lengths: Base refers to a model trained on data with random CoT lengths, while Opt refers to a model trained on data with optimal CoT lengths.
During testing, we allow the model to choose the CoT types on its own. The results shown in Figure 6 demonstrate that with a carefully designed training dataset, a smaller model can achieve significantly better performance—nearly 100% accuracy—even surpassing a larger model trained on base dataset.
However, in real-world datasets, the same problem may be associated with reasoning chains of varying lengths, none of which necessarily correspond to the optimal one. This experiment highlights the importance of carefully selecting CoT length when training models for chain-of-thought reasoning.
Algorithm 1 Length-filtered Vote
2:Output: Final Answer $\hat{a}$
4:Define $\mathcal{A}(c)$ as the corresponding answer of candidates $c$.
7:     for $a\in\mathcal{A}$ do
$p_{j}[a]=\frac{\sum_{c\in L_{j}}\mathbb{I}(\mathcal{A}(c)=a)}{|L_{j}|}$
8:     end for
9:end for
12:return $\hat{a}$
6 Length-filtered Vote
Section 5.2 highlights the importance of aligning the CoT length with the model’s capabilities and the task’s difficulty. However, achieving this alignment requires an accurate estimation of both the task and the model. Moreover, given a pretrained model, how can we leverage the optimal CoT length without any prior estimation of the task or even the model itself?
In this section, we propose a length-aware variant of majority vote, Length-filtered vote, where we use prediction uncertainty as a proxy to filter reliable CoT lengths. As in majority vote, given a model $f_{\theta}$, a question $q$, a ground truth answer $a^{*}$, we first sample a set of answer candidates independently $c_{1},\dots,c_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}f_{\theta}(q)$. After that, instead of direct vote, we group the answers based on their corresponding CoT length $\ell(c_{i})$ (discussed in Appendix A.1) into groups with equal bandwidth $D$ (by default, we set $D=2$), denoted as $\{L_{j}\}_{j=1}^{m}$. As our theory suggests that the prediction accuracy of CoT paths is peaked around a certain range of CoT length, we identify such groups through the prediction uncertainty of the answers within each group, based on the intuition that lower uncertainty implies better predictions. Specifically we calculate the Shannon entropy of the final answers given by the CoT chains in each group $L_{i}$, denoted as $H(L_{i})$. We then select $K$ (by default, we set $K=3$) out of $M$ groups that has the smallest entropy and then perform majority vote only on these selected groups. We summarize it in Algorithm 1.
Table 1: Performance of different models with varying sample numbers for Direct Vote on all candidates and Length-filtered Vote.
Model Method Number of Samples
20 30 40 50
Llama3.1-8B-Ins Direct Vote 35% 38% 39% 38%
Length-filtered Vote 36% 42% 42% 41%
Qwen2.5-7B-Ins Vote 34% 35% 36% 34%
Length-filtered Vote 36% 40% 38% 40%
We evaluate the propose method against vanilla majority vote (i.e., self-consistency) (Wang et al., 2023) on a randomly chosen subset of 100 questions from the GPQA dataset, a more challenging collection of multiple-choice questions. The results in Table 1 show that our filtered vote consistently outperforming vanilla majority vote at different sample numbers and have little performance degradation as the sample number increases. This indicates the importance of considering the influence CoT length in the reasoning process.
7 Conclusion
In this paper, we conduct experiments on a simplified synthetic dataset, drawing clear conclusions on how the CoT length affects the final performance. Our study also provides valuable insights, showing that the optimal CoT length should adapt to both model size and task difficulty. Furthermore, we present a rigorous theoretical framework demonstrating the non-monotonic scaling behavior of CoT length and how it is influenced by model size and task difficulty. Additionally, we conduct experiments on real-world datasets, yielding similar results. We also propose methods that can benefit from the optimal CoT length during both training and test phases. In this way, our analysis offers concrete theoretical and empirical insights into developing LLMs that adaptively select the appropriate reasoning length, avoiding either overthinking or underthinking.
Appendix

### v3

@@ 6 Related Work
Chain-of-Thought for LLM Reasoning. CoT has become a core technique for LLMs to solve complex reasoning tasks by generating intermediate steps (Wei et al., 2022). Numerous variants arise to enhance CoT reasoning with more structural substeps, such as least-to-most prompting (Zhou et al., 2023), tree of Thoughts (Yao et al., 2023), and divide-and-conquer methods (Zhang et al., 2024; Meng et al., 2024). These methods fundamentally treat CoT as a framework for task decomposition and subtask solving that falls in our analysis in Section 4.
Overthinking in CoT Reasoning. With the rise of powerful reasoning models like OpenAI o1, scaling test-time compute with long CoT has gained prominence (Snell et al., 2024; Chen et al., 2024d; Wu et al., 2024; Brown et al., 2024). These studies often suggest that more computation like longer CoT can lead to better results. However, this is not always true. With similar interests as ours, a few concurrent works also investigated the “overthinking” phenomenon (Chen et al., 2024c) where reasoning models generate excessively long CoTs for simple problems and proposed some mitigation strategies Han et al. (2024); Luo et al. (2025); Ma et al. (2025); Sui et al. (2025). Our analysis not only reveals the inverted-U curve of CoT length and the existence of optimal CoT length, but also provides a in-depth understanding on the scaling behaviors and simplicity bias of the optimal CoT length, as supported by both controlled experiments and theoretical analysis. This establishes a systematic explanation of overthinking and points out principled guidelines for better CoT designs.
Simplicity Bias and Occam’s Razor in Machine Learning. The simplicity bias of CoT identified in our work resonates with broader principles like Occam’s Razor, which favors simpler explanations or models. In machine learning, a ’simplicity bias’ often refers to neural networks learning simpler functions first or being biased towards solutions with lower intrinsic complexity (Arpit et al., 2017; Huh et al., 2021). Our findings extend this understanding to the realm of generated reasoning paths: we reveal that even structured, multi-step reasoning processes like CoT, as produced by LLMs, exhibit such simplicity bias by favoring concise reasoning paths, particularly as model capability increases.
Theoretical Understanding of CoT. Numerous studies aim to theoretically formalize the Chain-of-Thought (CoT) process and understand its effectiveness. They include analyzing CoT’s computational advantages via circuit complexity (Feng et al., 2023; Li et al., 2024b), demonstrating how coherent reasoning paths enhance error correction and accuracy (Cui et al., 2024), and quantifying step-wise information gain from an information-theoretic standpoint (Ton et al., 2024). Further research has shown that detailed CoT improves learning stability by affecting gradient dynamics (Li et al., 2024a), while controlled synthetic experiments have helped uncover underlying problem-solving mechanisms in LLMs (Ye et al., 2024). Distinct from these varied theoretical explorations, our theory characterizes how CoT length influences final performance and explains its scaling behaviors through the interplay of task decomposition and error accumulation. Furthermore, our findings on CoT scaling behaviors and the consequent need for model-specific CoT structures (as discussed in Section 3.2) resonate with the concept of algorithmic alignment (Xu et al., 2019), which suggests that models perform best when the problem structure aligns with their computation structure.
7 Conclusion
In this paper, we challenged the notion that longer Chain-of-Thought (CoT) processes are invariably superior, demonstrating through extensive experiments and theoretical analysis that CoT length and accuracy typically follow an inverted U-shaped curve, implying an optimal length that balances task decomposition against error accumulation. We discovered the simplicity bias of CoT, where more capable models prefer shorter effective reasoning paths, and formally derived scaling laws for this optimal length relative to model capability and task difficulty. Practically, we showed that reinforcement learning can guide models towards this optimal CoT length, that training on optimally-lengthed CoTs boosts performance, and proposed "Length-Filtered Vote" as a promising inference strategy. Our work underscores the critical need to calibrate CoT length, moving beyond a one-size-fits-all approach towards a principled framework where LLMs adaptively choose the right amount of thought to optimize reasoning.
Acknowledgement
Yisen Wang was supported by National Key R&D Program of China (2022ZD0160300), National Natural Science Foundation of China (92370129, 62376010), and Beijing Nova Program (20230484344, 20240484642). Yifei Wang and Stefanie Jegelka were supported in part by the NSF AI Institute TILOS (NSF CCF-2112665), and an Alexander von Humboldt Professorship.
Appendix
[…]
Appendix A Limitations
This proof-of-concept study highlights the critical role of aligning Chain-of-Thought (CoT) length with model capability and task difficulty, particularly in training data. However, accurately estimating the optimal CoT length in real-world scenarios remains challenging due to the complexity and variability of reasoning tasks. While our theoretical and empirical analyses offer practical heuristics for coarse approximation, they may not fully capture the nuances of diverse problem domains or model behaviors. We leave the development of more precise estimation methods and adaptive strategies for optimal CoT length selection in complex, real-world settings as promising directions for future work.
Appendix B Formal Definitions of Simplified Arithmetic Problem
@@
To begin, we aim to empirically investigate the relationship between reasoning performance and CoT length. Therefore, we need to control a given model to generate reasoning chains of varying lengths for a specific task. Unfortunately, no existing real-world dataset or model fully meets these strict requirements. Real-world reasoning tasks, such as GSM8K or MATH [Cobbe et al., 2021, Hendrycks et al., 2021b], do not provide multiple solution paths of different lengths, and manually constructing such variations is challenging. Moreover, it is difficult to enforce a real-world model to generate a diverse range of reasoning paths for a given question. Given these limitations, we begin our study with experiments on synthetic datasets.
B.1 Problem Formulation
To investigate the effect of CoT length in a controlled manner, we design a synthetic dataset of simplified arithmetic tasks with varying numbers of reasoning steps in the CoT solutions.
Definition B.1 (Problem).
@@ B.1 Problem Formulation
To ensure that CoT solutions of the same length have equal difficulty for a specific problem, we assume that each reasoning step performs the same operations within a single CoT process.
Definition B.2 (Solution).
@@
B.2 Contrast to vanilla arithmetic problem
Appendix C Supplementary Details on Real world Experiment for Optimal CoT Length
C.1 Implementation Details
Solution Length Control. To study the impact of CoT length on performance under a given problem difficulty, we need to induce the model to naturally generate solutions of varying lengths. Simply adding prompts like “please use 100 tokens to solve this problem” or “please use 10 steps to solve this problem” is not ideal because the model’s ability to follow instructions regarding output length is limited, and such fixed-length prompts may not ensure fairness across problems of different difficulties. Moreover, prompting for a specific length might lead the model to generate irrelevant tokens or steps just to “pad the length,” without actually changing the number of steps or the complexity of the reasoning. Additionally, controlling max_length is also problematic, as overly long responses might get truncated, which would directly lead to lower accuracy for longer outputs. What we really want is for the model to generate a complete and coherent long response on its own, so we can observe the corresponding accuracy.
To create solutions with varying step lengths with different complexity, we follow [Fu et al., 2023] by using in-context examples (8-shots) with three different levels of complexity to guide the model in generating solutions with different step counts. For each set of in-context examples, we sample 20 times, resulting in a total of 60 samples per question.
Step Segmentation. Simply measuring CoT length by counting tokens is neither rigorous nor meaningful. Since our focus is on final performance rather than efficiency, we care more about using CoT length to reflect the complexity of the reasoning pattern. In this sense, the number of reasoning steps can serve as a more appropriate indicator of CoT length. As we discussed earlier, the step number captures how the model decomposes the problem, which directly reflects the complexity of its reasoning. In contrast, token length fails to capture this because, as the model thinks more deeply and the number of steps increases, the number of tokens per step may decrease—making the total token count unpredictable and unreliable as a proxy for reasoning complexity.